determine sin theta to three decimal places when tan theta=1/2 and theta is an angle in standard position


1
Expert's answer
2021-10-06T16:38:34-0400

To find sin theta using the tan theta, firstly we should express the cot theta through the tan theta, using the trigonometric identity


"tan \u03b8*cot\u03b8=1;"

"cot\u03b8=1\/tan\u03b8=1\/1\/2=2;"

Then we can use another trigonometrical feature:


"sinθ=\sqrt{1/(1+cot^2θ)}=\sqrt{1/(1+2^2)}="

"=\sqrt{1/(1+4)}=\sqrt{1/5}=1/\sqrt{5}=0.447"